# Understanding and Mitigating Membership Inference Risks of Neural Ordinary Differential Equations

### v1

@@
Abstract.
Neural ordinary differential equations (NODEs) are an emerging paradigm in scientific computing for modeling dynamical systems. By accurately learning underlying dynamics in data in the form of differential equations, NODEs have been widely adopted in various domains, such as healthcare, finance, computer vision, and language modeling. However, there remains a limited understanding of the privacy implications of these fundamentally different models, particularly with regard to their membership inference risks.
@@ Abstract.
In this work, we study the membership inference risks associated with NODEs. We first comprehensively evaluate NODEs against membership inference attacks. We show that NODEs are twice as resistant to these privacy attacks compared to conventional feedforward models such as ResNets. By analyzing the variance in membership risks across different NODE models, we identify the factors that contribute to their lower risks. We then demonstrate, both theoretically and empirically, that membership inference risks can be further mitigated by utilizing a stochastic variant of NODEs: Neural stochastic differential equations (NSDEs). We show that NSDEs are differentially-private (DP) learners that provide the same provable privacy guarantees as DP-SGD, the de-facto mechanism for training private models. NSDEs are also effective in mitigating existing membership inference attacks, demonstrating risks comparable to private models trained with DP-SGD while offering an improved privacy-utility trade-off. Moreover, we propose a drop-in-replacement strategy that efficiently integrates NSDEs into conventional feedforward models to enhance their privacy.
Neural ODEs, Differential privacy, Membership inference
copyright: nonejournalyear: 2025copyright: rightsretainedconference: the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwandoi: xxxxxxx.xxxxxxxisbn: 978-1-4503-xxxx-x/18/06
1. Introduction
In recent years, neural ordinary differential equations (NODEs) (Chen et al., 2018) have been gaining traction across various domains. This new family of neural networks enables modeling of underlying data distributions through learnable differential equations—a fundamental tool for describing processes in the physical world. By shifting from manually designing governing equations to data-driven modeling, NODEs have been deployed in various fields, such as time-series forecasting (Chen et al., 2018; Kidger et al., 2020; Morrill et al., 2021; Rubanova et al., 2019; De Brouwer et al., 2019; Jeon et al., 2022; Wu et al., 2024), scientific machine learning (ML) (Greydanus et al., 2019; Zhong et al., 2020; Rackauckas et al., 2020; Verma et al., 2024; Lee et al., 2021; Gruber et al., 2024), or general ML tasks (Chen et al., 2018; Dupont et al., 2019a; Yildiz et al., 2019; Xia et al., 2021; Song et al., 2021; Cho et al., 2022, 2024). Moreover, applications of NODEs in finance or healthcare (Rubanova et al., 2019; Irie et al., 2022; Lee et al., 2024), which heavily rely on time-series or physics-informed modeling, are becoming closely integrated into our daily lives.
Since NODEs are deployed in fields where the training data often contains confidential or sensitive information, such as patient medical records from hospitals, this naturally raises privacy concerns. Recent work has shown the risk of membership inference (Shokri et al., 2017), where an adversary can determine if a specific data record was part of the training set. Membership inference has been extensively studied in the context of feedforward neural networks, such as convolutional neural networks (Yeom et al., 2018; Shokri et al., 2017; Song and Mittal, 2021; Carlini et al., 2022) or Transformer-based language models (Zhang et al., 2024). However, this risk has not been studied against NODEs that operate differently, with their internal structures designed based on principles from mathematical modeling. It is unclear whether applications using these models face similar privacy risks, and how to mitigate them while preserving usability.
In this work, we study the privacy implications of employing differential equation-based neural networks, such as NODEs. Specifically, we address the following four research questions:
Q1
@@ 1. Introduction
How vulnerable are NODEs to membership inference attacks? We address this question empirically through a systematic evaluation. We implement a risk analysis framework that runs five off-the-shelf membership inference attacks across a variety of neural networks including NODEs and baseline feedforward models. In evaluation on two datasets (CIFAR-10 and CIFAR-100), commonly used in prior work on assessing membership risks, we demonstrate that NODEs exhibit up to 2×2$\times$2 × lower membership risks compared to feedforward neural networks, like ResNets (He et al., 2016). We also find that this is partly due to NODEs exhibiting less overfitting while achieving accuracy comparable to conventional neural network models.
@@
Moreover, we analyze factors that may contribute to this membership risks. We first examine configurations one can control at the block-level: solvers, tolerance, and step size. Our analysis indicates that these factors, which affect the complexity of learned dynamics while keeping the same model capacity, can increase/decrease the likelihood of overfitting. We also test the model-level configurations: the number of blocks and advanced variants of NODEs (Dupont et al., 2019a; Xia et al., 2021; Norcliffe et al., 2020). They increase the model capacity or the complexity of learned dynamics, and therefore, increase membership risks.
Q2
@@
How can we mitigate the membership risks of NODEs? Our previous evaluation shows that while NODEs are less vulnerable, the risks to membership inference still exist. We next address this problem by focusing on a stochastic adaptation of NODEs: neural stochastic differential equations (NSDEs) (Liu et al., 2019). NSDEs, similar to NODEs, model the underlying dynamics with ODEs but augment them with a diffusion term that incorporates Gaussian noise.
@@
We first formally show that this diffusion term functions as a differentially-private (DP) mechanism during training, and therefore, NSDEs become DP-learners. We establish a theoretical bound on the privacy leakage of NSDEs at each iteration of mini-batch SGD and then present a mechanism for accounting for the total privacy leakage during training. Surprisingly, we show that the total privacy leakage ($\varepsilon$) guaranteed by the NSDE training process is the same as that of the de-facto standard for training private models, DP-SGD (Abadi et al., 2016). The leakage $\varepsilon$ is proportional to the number of training iterations and the level of the Gaussian noise ($\sigma$).
@@
We then empirically evaluate the effectiveness of NSDEs in further reducing the membership inference risks. We demonstrate that NSDEs can make existing membership inference attacks 1.8–10$\times$ less effective compared to their effectiveness on NODEs. We observe that the utility of those NSDE models, under the same privacy guarantees, is comparable to (or slightly higher than) that of ResNet models trained with DP-SGD. This shows their potential as a countermeasure with an improved privacy-utility trade-off. We also compare NSDEs with heuristic defenses (Kaya et al., 2020) designed to counter membership inference attacks but lacking provable privacy guarantees. We show that NSDEs are 22225×5$\times$5 × more effective in mitigating the membership inference risks.
Q3
@@
How can existing pre-trained models be augmented with NSDEs to enhance privacy? Implementing our privacy mechanism requires training a new NSDE model from scratch, which can be a computationally demanding option for service providers. We address this practical challenge by efficiently applying NSDEs to existing pre-trained models, particularly conventional feedforward models like ResNets. We leverage the common practice of transfer learning in supervised learning, where the last layer of a pre-trained model are replaced with a classification head, followed by fine-tuning on data over a few iterations. Our strategy, replace-then-finetune, replaces the last few layers of a pre-trained model with an NSDE block. We test our strategy on ResNet14 pre-trained on CIFAR-10. Our evaluation shows that fine-tuned models achieve accuracy comparable to the pre-trained models, while reducing the membership risks by 10$\times$ in CIFAR-10. We hope our work encourages the adoption of NSDEs as a viable mechanism for further reducing privacy risks while maintaining model utility.
@@
  • We are the first to study the membership inference risks of NODEs, which utilize a fundamentally different approach to modeling data distributions. Our comprehensive evaluation shows that NODEs exhibit 2$\times$ lower membership risks compared to feedforward neural networks like ResNets. We also analyze the factors contributing to this reduced risk.
  • We propose NSDEs as a defensive mechanism for mitigating the membership inference risks of NODEs. We formally show that NSDEs are DP learners, offering the same provable privacy guarantees as DP-SGD. We also present an accounting mechanism to track the total privacy leakage during NSDE training.
  • We show, in our empirical evaluation that NSDEs achieve a 2–5$\times$ reduction in membership risks compared to defenses without provable guarantees applied to NODEs. Compared to private models trained with DP-SGD, NSDEs demonstrate the same effectiveness in mitigating membership inference attacks while achieving slightly higher accuracy.
  • We present a practical strategy to leverage this privacy benefit without fully replacing existing models with NSDEs, replace-then-finetune. It reduces membership risks by 10$\times$ while achieving a higher utility than the models trained with DP-SGD.
2. Background
2.1. Neural Ordinary Differential Equations
@@
where ${}h(t)$ are the state variables and ${}f_{\Theta}$ is a neural network, parameterized by its parameters $\Theta$, defining the rate of change in ${}h(t)$. The scalar variable $t$ typically refers to a time variable and the system states correspond to observables in time-dependent physical processes (Strogatz, 2018) or measurements in time-series data (e.g., (Rubanova et al., 2019)). However, in the context of hidden representations within a deep feedforward network (FFN), $t$ instead refers to a continuous depth parameter. More precisely, recall that the flow map ${}\varphi=~{}\varphi(t,~{}h)$ of an ODE system associates to each state ${}h$ its “flowed out” state after evolving along the ODE for $t$ time, i.e., ${}\varphi(0,~{}h)=~{}h$ and $(d/dt)~{}\varphi(t,~{}h)=~{}f_{\Theta}(~{}h(t),t)$ in the case of the NODE above. Then, the number of “layers” in the NODE architecture defined by ${}f_{\Theta}$ is analogous to the number of “time steps” used to generate the flowed-out state ${}h(t)$, and hence each network prediction based on this quantity. Our focus is on this use case, which allows for depth control through direct manipulation of the ODE time-stepping.
@@
Different choices of numerical time-integration yield different realizations of FNNs, e.g., under the forward Euler method with step size 1, autonomous ($t$-independent) NODEs degenerate to ResNets (He et al., 2016), ${}h(t_{i+1})=~{}h(t_{i})+~{}f_{\Theta}(~{}h(t_{i}))$. With this setup, the forward pass for both training and inference is equivalent to recurrently applying a nonlinear transformation, defined by ${}f_{\Theta}$, to a hidden state, followed by a skip connection. In classification tasks, the last hidden representation ${}h(t_{n})$ is fed into a classifier network, producing logits, and the loss of the model is obtained via computing the cross-entropy. The continuous-depth network ${}M(t,~{}h)=~{}M_{t}~{}(~{}h)$ can then be thought of as discretizing the ODE flow map, i.e., ${}M_{t_{n}}(~{}h_{0})=~{}h(t_{n})$, where ${}h(t_{n})$ is the solution of the IVP defined above. As mentioned before, any intermediate time steps $t_{1},...,t_{n-1}$ used in the production of the final state ${}h(t_{n})$ are analogous to the labels on hidden layers in an FFN.
For the backward process required during training, NODEs use either standard backpropagation through the rolled-out computational path, or the adjoint sensitivity method (Pontryagin, 2018). The primary challenge lies in achieving stable and memory-efficient gradient computation for ${}f_{\Theta}$ with respect to its parameters, as backpropagating directly through an iterative numerical solver can be memory intensive and exacerbate vanishing or exploding gradient issues.
NODE variants. Many variants have been developed to improve their expressivity and efficiency. Augmented NODEs (ANODEs) (Dupont et al., 2019b) propose an additional dimension to the state variable $[~{}h(t),~{}a(t)]$, to increase expressivity. ANODEs were further extended into second-order NODEs (SONODEs) (Norcliffe et al., 2020), which formulates a system of coupled first-order NODEs, implementing the broader class of augmented models. Other variants incorporate concepts from gradient descent optimization, such as momentum, resulting in heavy ball NODEs (HBNODEs) (Xia et al., 2021) and Nesterov NODEs (Nguyen et al., 2022), both of which improve computational efficiency by reducing the number of function evaluations in the forward pass.
NSDEs (Liu et al., 2019; Li et al., 2020; Kidger et al., 2021a) extend NODEs to the modeling of stochastic processes, typically introducing randomness in the form of additive Gaussian noise. These systems can be formulated in terms of deterministic, vector-valued “drift” and stochastic, matrix-valued “diffusion” terms as follows:
@@
where $d~{}B_{t}$ denotes the Wiener increment—a zero mean, unit variance, delta-correlated Gaussian process describing “white noise”. A simple but popular discretization of NSDEs is given by the Euler–Maruyama method (Platen and Bruti-Liberati, 2010) with step size 1, expressed as ${}h(t_{i+1})=~{}h(t_{i})+~{}f_{\Theta_{1}}(~{}h,t_{i})\Delta t+~{}G_{\Theta_{2}}(~{}h,t_{i})~{}w_{t_{i}}$, where $\Delta t$ is the time step size and ${}w_{t_{i}}\sim\mathcal{N}(0,\Delta t)$ is randomly sampled. In this formulation, Gaussian noise is injected alongside each residual connection. Because of this property, NSDEs have shown enhanced robustness against adversarial examples (Liu et al., 2020) compared to standard NODEs. Our study further explores this property as a formal mechanism for mitigating privacy risks.
2.2. Membership Inference Attacks
Membership inference attacks aim to determine whether a specific example is a member of the training data. To do this, the attacker exploits the difference in the target model’s response to the specific example when it is a member versus a non-member. Membership inference can be considered a threat on its own, but a model’s vulnerability to inference attacks also implies its potential to leak other private information outside of this context, an idea which aligns closely with the definition of differential privacy (Dwork, 2006) as a worst-case certification.

### v2

# Modeling Neural Networks with Privacy Using Neural Stochastic Differential Equations

@@
Abstract
@@
In this work, we study the feasibility of using neural ordinary differential equations (NODEs) to model systems with intrinsic privacy properties. Unlike conventional feedforward neural networks, which have unlimited expressivity and can represent arbitrary mappings between inputs and outputs, NODEs constrain their learning to the solution of a system of differential equations. We first examine whether this constraint reduces memorization and, consequently, the membership inference risks associated with NODEs. We conduct a comprehensive evaluation of NODEs under membership inference attacks and show that they exhibit twice the resistance compared to conventional models such as ResNets. By analyzing the variance in membership risks across different NODE models, we find that their limited expressivity leads to reduced overfitting to the training data. We then demonstrate, both theoretically and empirically, that membership inference risks can be further mitigated by utilizing a stochastic variant of NODEs: neural stochastic differential equations (NSDEs). We show that NSDEs are differentially-private (DP) learners that provide the same provable privacy guarantees as DP-SGD, the de-facto mechanism for training private models. NSDEs are also effective in mitigating membership inference attacks, achieving risk levels comparable to private models trained with DP-SGD while offering an improved privacy-utility trade-off. Moreover, we propose a drop-in-replacement strategy that efficiently integrates NSDEs into conventional feedforward architectures to enhance their privacy.
1 Introduction
Neural networks are “universal function approximators”: they can learn arbitrary continuous mappings from training data to target outputs [1]. This expressivity allows highly parameterized networks to model complex, non-linear relationships and capture hidden patterns that traditional methods may struggle to detect. However, this strength also introduces an undesirable property: memorization [2]. Memorization occurs when the learned mappings are overly specific to individual training instances—regardless of whether these mappings generalize to unseen data—potentially leading to the leakage of their membership through model queries [3, 4, 5, 6, 7]. Given that neural networks are often deployed in sensitive domains, such as healthcare, where training data may include confidential medical records, this membership leakage naturally raises privacy concerns.
Recent work has introduced a new class of neural networks: neural ordinary differential equations (NODEs) [8]. Unlike conventional neural networks, NODEs constrain the learned mappings to solutions of parameterized differential equations—a fundamental mathematical tool for describing processes in the physical world. By shifting from manually designing governing equations to data-driven modeling, NODEs have emerged as a powerful paradigm in scientific computing for modeling dynamical systems across various domains, such as time-series forecasting [9, 10, 11], scientific machine learning (ML) [12, 13, 14, 15], and computer vision tasks [16, 17, 18, 19]. These works demonstrate that NODEs, despite their constrained expressivity to learnable differential equations, can achieve performance comparable to that of conventional neural networks across a range of tasks.
In this work, we study the privacy implications of employing differential equation-based neural networks. Specifically, we ask the following research question: Does this emerging class of neural networks (and their variants) offer inherent privacy advantages, such as reduced membership inference risks, compared to conventional neural networks? Because NODEs restrict model expressivity to the system of differential equations, they may influence memorization and overfitting—key factors attributing to membership inference risks. Their stochastic variants, neural stochastic differential equations (NSDEs), introduce randomness into the model dynamics and thus, may satisfy formal privacy guarantees. Moreover, the modular design of these models could enable their integration into conventional feedforward architectures as privacy-enhancing components.
@@ 1 Introduction
We first conduct a systematic evaluation of the membership inference risks associated with NODEs. To this end, we develop a risk analysis framework that runs six off-the-shelf membership inference attacks across a variety of neural networks including NODEs and baseline feedforward models. In our evaluation on four image recognition benchmarks, commonly used in prior work on assessing membership risks, we find that NODEs exhibit up to 2×2$\times$2 × lower membership risks compared to conventional feedforward networks, like ResNets [20]. We also find that this reduced risk is partly attributable to the lower degree of overfitting in NODEs, while achieving accuracy comparable to baselines.
@@
Moreover, we analyze factors that may affect their membership risks. We first examine configurations one can control at the block-level: solvers, tolerance, and step size. Our analysis indicates that these factors, which change the complexity of learned dynamics while keeping the same model capacity, can increase/decrease the likelihood of overfitting. We also test the model-level configurations: the number of blocks and advanced variants of NODEs [16, 18, 21]. They increase the model capacity or the complexity of learned dynamics, and therefore, increase membership risks.
@@
Our previous evaluation shows that, while NODEs are less vulnerable, the risks to membership inference still exist. We address this problem by employing a stochastic adaptation of NODEs: NSDEs [22]. NSDEs also model underlying dynamics using differential equations, but augment them with a diffusion term that introduces Gaussian noise.
@@
We first formally show that this diffusion term acts as a differentially-private (DP) mechanism during training, and turns NSDEs into DP-learners. We establish a theoretical bound on privacy leakage per mini-batch SGD iteration and present a mechanism to account for total privacy leakage over the entire training process. Surprisingly, we find that the total privacy leakage ($\varepsilon$) guaranteed by NSDE training matches that of DP-SGD [23], the de-facto standard for training private models. The leakage $\varepsilon$ is proportional to the number of training iterations and the noise level ($\sigma$).
@@
We then empirically evaluate the effectiveness of NSDEs in further reducing the membership inference risks. Our results show that NSDEs reduce the effectiveness of existing membership inference attacks by 1.8–10$\times$ compared to NODEs. Despite offering the same privacy guarantee, NSDE models maintain utility that is comparable to (or slightly higher than) ResNet models trained with DP-SGD. This shows NSDEs as a promising countermeasure that achieves an improved privacy-utility trade-off. We also compare NSDEs with prior heuristic defenses [24] designed to counter membership inference attacks but lacking formal privacy guarantees. We show that NSDEs are 22225×5$\times$5 × more effective in mitigating the membership inference risks.
@@
Implementing our privacy mechanism requires training an NSDE model from scratch, which can be a computationally demanding option for model providers. To address this practical challenge, we propose an effective strategy to apply NSDEs to existing pre-trained models, particularly conventional model architectures like ResNets. We leverage the common practice of transfer learning where the last layer of a pre-trained model is replaced with a classification head, followed by fine-tuning on data for a few iterations. Our proposed strategy, replace-then-finetune, substitutes the last few layers of a pre-trained model with an NSDE block. We test our strategy on ResNet14 pre-trained on the four benchmark tasks. Across all benchmarks, the fine-tuned models maintain accuracy comparable to the pre-trained models, while reducing the membership risks substantially (e.g., by 10$\times$ in CIFAR-10). We hope our work encourages the adoption of NSDEs as a viable mechanism for further reducing privacy risks while maintaining model utility.
@@
  • We are the first to study the membership inference risks of NODEs. Our comprehensive evaluation shows that NODEs exhibit 2$\times$ lower membership risks compared to baseline models. We attribute this to NODEs’ reduced overfitting, which is from their restricted expressivity—learning within the constraints of a system of ODEs.
  • We propose NSDEs as a defense mechanism to mitigate membership inference attacks. We formally show that NSDEs provide the same provable privacy guarantee as DP-SGD. We also present an accounting mechanism to track the total privacy leakage during NSDE training.
  • In our evaluation, we show that NSDEs reduce membership inference risks by 2–5$\times$ compared to heuristic defenses without provable guarantees. Compared to private models trained with DP-SGD, NSDEs offer comparable (provable) protection against membership inference attacks while achieving slightly higher accuracy.
  • We present a practical strategy to leverage these privacy benefits without fully replacing existing models with NSDEs: replace-then-finetune. This strategy reduces membership inference risks by 10$\times$ while achieving a higher utility than private models traind with DP-SGD.
2 Background
2.1 Neural Ordinary Differential Equations
@@
where ${}h(t)$ are the state variables and ${}f_{\Theta}$ is a neural network, parameterized by its parameters $\Theta$, defining the rate of change in ${}h(t)$. The scalar variable $t$ refers to a time variable and the system states correspond to observables in time-dependent physical processes [25] or measurements in time-series data (e.g., [26]). However, in the context of hidden representations within a deep feedforward network (FFN), $t$ instead refers to a continuous depth parameter. More precisely, recall that the flow map ${}\varphi=~{}\varphi(t,~{}h)$ of an ODE system associates to each state ${}h$ its “flowed out” state after evolving along the ODE for $t$ time, i.e., ${}\varphi(0,~{}h)=~{}h$ and $(d/dt)~{}\varphi(t,~{}h)=~{}f_{\Theta}(~{}h(t),t)$ in the case of the NODE above. Then, the number of “layers” in the NODE architecture defined by ${}f_{\Theta}$ is analogous to the number of “time steps” used to generate the flowed-out state ${}h(t)$, and hence each network prediction based on this quantity. Our focus is on this use case, which allows for depth control through direct manipulation of the ODE time-stepping.
@@
Different choices of numerical time-integration yield different realizations of FNNs, e.g., under the forward Euler method with step size 1, autonomous ($t$-independent) NODEs degenerate to ResNets [20], ${}h(t_{i+1})=~{}h(t_{i})+~{}f_{\Theta}(~{}h(t_{i}))$. With this setup, the forward pass for both training and inference is equivalent to recurrently applying a nonlinear transformation, defined by ${}f_{\Theta}$, to a hidden state, followed by a skip connection. In classification tasks, the last hidden representation ${}h(t_{n})$ is fed into a classifier network, producing logits, and the loss of the model is obtained via computing the cross-entropy. The continuous-depth network ${}M(t,~{}h)=~{}M_{t}~{}(~{}h)\approx~{}\varphi(t,~{}h)$ can then be thought of as discretizing the ODE flow map, i.e., ${}M_{t_{n}}(~{}h_{0})=~{}h(t_{n})$, where ${}h(t_{n})$ is the solution of the IVP defined above. As mentioned before, any intermediate time steps $t_{1},...,t_{n-1}$ used in the production of the final state ${}h(t_{n})$ are analogous to the labels on hidden layers in an FFN.
For the backward process required during training, NODEs use either standard backpropagation through the rolled-out computational path, or the adjoint sensitivity method [27]. The primary challenge lies in achieving stable and memory-efficient gradient computation for ${}f_{\Theta}$ with respect to its parameters, as backpropagating directly through an iterative numerical solver can be memory intensive and exacerbate vanishing or exploding gradient issues.
NODE variants. Many variants have been developed to improve their expressivity and efficiency. Augmented NODEs (ANODEs) [16] propose an additional dimension to the state variable $[~{}h(t),~{}a(t)]$, to increase expressivity. ANODEs were further extended into second-order NODEs (SONODEs) [21], which formulates a system of coupled first-order NODEs, implementing the broader class of augmented models. Other variants incorporate concepts from gradient descent optimization, such as momentum, resulting in heavy ball NODEs (HBNODEs) [18] and Nesterov NODEs [28], both of which improve computational efficiency by reducing the number of function evaluations in the forward pass.
NSDEs [22, 29, 30] extend NODEs to the modeling of stochastic processes, typically introducing randomness in the form of additive Gaussian noise. These systems can be formulated in terms of deterministic, vector-valued “drift” and stochastic, matrix-valued “diffusion” terms as follows:
@@
where $d~{}B_{t}$ denotes the Wiener increment—a zero mean, unit variance, delta-correlated Gaussian process describing “white noise”. A simple but popular discretization of NSDEs is given by the Euler–Maruyama method [31] with step size 1, expressed as ${}h(t_{i+1})=~{}h(t_{i})+~{}f_{\Theta_{1}}(~{}h,t_{i})\Delta t+~{}G_{\Theta_{2}}(~{}h,t_{i})~{}w_{t_{i}}$, where $\Delta t$ is the time step size and ${}w_{t_{i}}\sim\mathcal{N}(0,\Delta t)$ is randomly sampled. In this formulation, Gaussian noise is injected alongside each residual connection. Because of this property, NSDEs have shown enhanced robustness against adversarial examples [32] compared to standard NODEs. Our study further explores this property as a formal mechanism for mitigating privacy risks.
2.2 Membership Inference Attacks
Membership inference attacks aim to determine whether a specific sample is a member of the training data. To do this, the attacker exploits the difference in the target model’s response to the specific sample when it is a member versus a non-member. Membership inference can be considered a threat on its own, but a model’s vulnerability to inference attacks also implies its potential to leak other private information outside of this context, an idea which aligns closely with the definition of differential privacy [33].